# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ VI-A Specific Phase of PPFL Model
[…]
@@
  • Weighted aggregation. The server selects some clients and aggregates their local gradients as a global gradient. Denote the selected client set as $S_{benign}$ with the size $m=|S_{benign}|$. The server aggregates all the $m$ ciphertexts $\{\overline{ct}_{id_{i}}\}_{{id_{i}}\in S_{benign}}$ as an aggregation ciphertext $\overline{ct}_{add}^{m}$. Specifically, the server first computes
    $\begin{split}\overline{ct}_{add}^{1}\leftarrow\texttt{SMHE.Add}&(\overline{ct}_{id_{1}},(0)_{m+1};id_{1},id_{m};\{cz_{id_{1}},\\ &\bm{\Gamma}_{id_{1}}\},\{cz_{id_{m}},\bm{\Gamma}_{id_{m}}\})\end{split}$
    For $i\in[2,m]$, the server computes
    $\begin{split}\overline{ct}_{add}^{i}\leftarrow\texttt{SMHE.Add}&(\overline{ct}_{add}^{i-1},\overline{ct}_{id_{i}};id_{i-1},id_{i};\{cz_{id_{i-1}},\\ &\bm{\Gamma}_{id_{i-1}}\},\{cz_{id_{i}},\bm{\Gamma}_{id_{i}}\}).\end{split}$
    Here, we have $id_{i-1}=id_{m}$ when $i=1$, and $id_{i+1}=id_{1}$ when $i=m$. The aggregation ciphertext is $\overline{ct}_{add}^{m}\overset{\triangle}{=}(\bar{c}_{0},\bar{c}_{id_{1}},\cdots,\bar{c}_{id_{m}})$ and $\bar{c}_{id_{i}}$ is then sent to the client $id_{i}$.
Partial decryption (Client side): Each client $id_{i}\ (id_{i}\in S_{benign})$ computes the partial decryption result $\nu_{id_{i}}=\texttt{SMHE.PartDec}(\bar{c}_{id_{i}},s_{id_{i}})=\bar{c}_{id_{i}}\cdot s_{id_{i}}+e_{id_{i}}\pmod{Q}$ with $e_{id_{i}}\leftarrow\chi$. Then, it sends $\nu_{id_{i}}$ to the server.
Full decryption and model update (Server side):
  • Full decryption. After receiving $\{\nu_{id_{i}}\}_{id_{i}\in S_{benign}}$ from all the $m$ clients in $S_{benign}$, the server obtains the final decryption result $\mu=\bar{c}_{0}+\sum_{id_{i}\in S_{bengin}}\nu_{id_{i}}\ (\bmod Q)$. Then the server decodes $\mu$ as the weighted aggregation gradient $\hat{g}_{t}$.
VI-B Security and Correctness
@@
The security of the PPFL model is guaranteed by the underlying security of the SMHE scheme, as analyzed in Section V-B1. We now analyze the correctness of the ciphertext aggregation and decryption processes within PPFL.
@@
During the aggregation process, the aggregated ciphertext is computed as $\overline{ct}_{\text{add}}^{1}\leftarrow\texttt{SMHE.Add}(\overline{ct}_{id_{1}},(0)_{m+1};id_{1},id_{m};$$\{cz_{id_{1}},\bm{\Gamma}_{id_{1}}\},\{cz_{id_{m}},\bm{\Gamma}_{id_{m}}\}),$ yielding:
$\overline{ct}_{\text{add}}^{1}=$ $(c_{0}^{id_{1}}+x_{0}^{id_{1},id_{m}}+z_{0}^{id_{m}}+x_{0}^{id_{m},id_{1}}+z_{0}^{id_{1}},$
$c_{1}^{id_{1}}+x_{1}^{id_{1},id_{m}}+z_{1}^{m},0,\ldots,0,x_{1}^{id_{m},id_{1}}+z_{1}^{id_{1}}),$
@@
$\bar{c}_{0}=\sum_{id_{i}\in S_{benign}}\left(c_{0}^{id_{i}}+x_{0}^{id_{i},id_{i-1}}+x_{0}^{id_{i},id_{i+1}}+2z_{0}^{id_{i}}\right),$
$\bar{c}_{id_{i}}=c_{1}^{id_{i}}+x_{1}^{id_{i},id_{i-1}}+x_{1}^{id_{i},id_{i+1}}+z_{1}^{id_{i-1}}+z_{1}^{id_{i+1}},i\in[1,m].$
Here, $m=|S_{benign}|$ and $cx_{id_{i},id_{j}}=(x_{0}^{id_{i},id_{j}},x_{1}^{id_{i},id_{j}})\leftarrow\texttt{Extend}(\bm{\Gamma}_{id_{i}},pk_{id_{i}},pk_{id_{j}})$ for $id_{i},id_{j}\in S_{benign}$, with $id_{i-1}=id_{m}$ when $i=1$, and $id_{i+1}=id_{1}$ when $i=m$.
@@
$\quad\mu=\bar{c}_{0}+\sum_{id_{i}\in S_{benign}}(\bar{c}_{id_{i}}\cdot s_{id_{i}}+e_{id_{i}})$ (12)
$=\sum_{id_{i}\in S_{benign}}\Big{(}\big{(}c_{0}^{id_{i}}+c_{1}^{id_{i}}s_{id_{i}}+e_{id_{i}}\big{)}+\langle sk_{id_{i}},cx_{{id_{i}},id_{i-1}}\rangle$
$\quad+\langle sk_{id_{i}},cz_{id_{i-1}}\rangle+\langle sk_{id_{i}},cx_{id_{i},id_{i+1}}\rangle+\langle sk_{id_{i}},cz_{id_{i+1}}\rangle\Big{)}$
$=\sum_{id_{i}\in S_{benign}}\big{(}c_{0}^{id_{i}}+c_{1}^{id_{i}}s_{id_{i}}\big{)}+e^{*}\approx\sum_{id_{i}\in S_{benign}}\mu_{id_{i}}\pmod{Q}.$
Eq. (12) holds due to the correctness of the masking scheme, which ensures that
@@
and similarly, $\langle sk_{id_{i}},cx_{{id_{i}},id_{i+1}}\rangle+\langle sk_{id_{i}},cz_{id_{i+1}}\rangle\approx 0\pmod{Q}.$ The error bound of the decryption is $\|e^{*}\|_{\infty}\leq(3m^{3}+6m^{2}+m)B_{\chi}+2dm^{2}\cdot B_{\chi}B_{\mathcal{H}}.$
VI-C Implementation
VI-C1 Experimental Setup
The experiments were conducted on a computer equipped with an Nvidia GeForce GTX 1080 Ti GPU and an Intel Core i7-6700 CPU. The SMHE algorithms were implemented in C++ using the NTL 10.4.0 [31] and GMP 6.2.1 [32] libraries to handle arbitrary-length integers and high-precision arithmetic. The FL models were executed using the PyTorch 1.11.0 framework in Python 3.8. To build our PPFL framework, a dynamic library containing all SMHE-related code was generated and invoked within a Python script. The code is available at https://github.com/JiahuiWu2022/SMHE.git.
Benchmark Models, Dateset, and Parameter Setting. To evaluate the effectiveness of our SMHE-based PPFL model, we conduct comparative experiments against the existing multiparty HE schemes: CDKS [21], THE [14], and MKGSW [17]. All models are assessed under two distinct learning scenarios to ensure comprehensive and fair comparison. In the first scenario, we train a fully connected neural network (FCN) on the MNIST dataset. The network includes an input layer with $784$ neurons, two hidden layers with $128$ and $64$ neurons respectively, and an output layer with $10$ neurons representing digit classes from 00 to $9$ using one-hot encoding. The model is optimized using the Adam algorithm with a mini-batch size of $64$. In the second scenario, we adopt the CIFAR-10 dataset to train a lightweight version of AlexNet with approximately $1.25$ million parameters. Training is conducted using a batch size of $128$ and the RMSprop optimizer with a decay factor of $10^{-6}$. For both FL tasks, the datasets are randomly partitioned among multiple clients for simulating collaborative training.
@@ VI-C1 Experimental Setup
VI-C2 Experimental Results
@@
Runtime. Fig. 3 presents the runtime breakdown of PPFL models across different computation phases on both the client and server sides for a single training iteration. The results are averaged over $100$ training iterations for all models except the MKGSW-based model, whose runtime is approximated using the MKGSW evaluation time reported in TFHE [35]. As a bit-level HE scheme, MKGSW incurs an impractically high runtime due to its inherent computational complexity. For example, under the FCN setting, the total runtime reaches $3570.65$s on the client side and $2702.17$s on the server side. This excessive overhead renders MKGSW unsuitable for real-world PPFL applications. In contrast, the other three models, THE-based, CDKS-based, and our SMHE-based models, exhibit significantly lower runtimes, making them more feasible for practical deployment. Among these, THE-based and CDKS-based models demonstrate comparable runtime performance. Our SMHE-based scheme introduces a moderate increase in runtime relative to THE and CDKS, primarily due to the incorporation of the masking extension mechanism. For instance, under the AlexNet setting, the total client-side runtime of SMHE increases by less than 2×2$\times$2 × compared to the CDKS- and THE-based models ($4.64$s vs. $2.72$s and $7.77$s, respectively). Similarly, on the server side, SMHE incurs a higher overall runtime – $2.17$s for FCN and $38.22$s for AlexNet – compared to CDKS ($0.97$s and $31.23$s) and THE ($0.96$s and $32.58$s). In all cases, the runtime increase remains within a 2×2$\times$2 × range, which we consider a reasonable trade-off given the enhanced security guarantees provided by SMHE.
Besides, compared to THE, our approach significantly reduces the client-side key generation time (e.g., $0.08$s vs. $5.09$s for FCN) by allowing each participant to generate their key locally, without the need for interactive threshold key generation protocols. This eliminates the communication overhead inherent in THE’s distributed key setup, leading to a substantial reduction in key generation and distribution latency.
Overall, SMHE strikes a practical balance between improved security, flexibility, and runtime efficiency. Compared to MKGSW, it achieves vastly lower computational overhead, while maintaining acceptable runtime increases relative to THE and CDKS. These features make SMHE a scalable and effective solution for PPFL applications.
TABLE II: Comparison of communication traffic for one iteration of different PPFL models.
@@
Communication Traffic. Table II presents the communication traffic per iteration for different PPFL models under FCN and AlexNet networks. The row labeled “Plain (#gradient)” reflects the number of model gradients in the plain FL, $109{,}386$K for FCN and $1.25$M for AlexNet. For encrypted models, the MKGSW-based model introduces an extremely high communication burden, $33.39$GB per iteration for FCN and $381.47$GB for AlexNet, mainly due to its bit-level encryption. This renders MKGSW impractical for real-world PPFL applications, where communication efficiency is a critical concern. In contrast, the other three models: THE-based, CDKS-based, and our SMHE-based scheme offer significantly lower communication costs and are more suitable for practical deployment. Specifically, THE-based and CDKS-based models share similar ciphertext formats and therefore exhibit identical communication traffic, requiring only $7$MB and $76.5$MB per iteration for FCN and AlexNet, respectively. Our SMHE-based model incurs a moderate increase in communication traffic, reaching $16$MB for FCN and $139.5$MB for AlexNet. This overhead stems from the ciphertext extension mechanism introduced by our masking design to support secure multi-key computation. However, despite the increase, the total traffic remains within approximately 2×2$\times$2 × that of CDKS and THE, which is a reasonable cost given the added benefits. Overall, SMHE offers a practical and scalable solution that balances enhanced security and key flexibility with acceptable communication overhead, unlike MKGSW, whose bit-level structure incurs prohibitive costs in PPFL settings.
@@
Model Accuracy. Fig. 4 displays the final model accuracy of the PPFL models under different client elimination rates. As observed from the results, the SMHE (CDKS)-based model exhibits strong robustness against client elimination. Even with $50$ distributed clients and a high elimination rate of $50\%$, the final model accuracy remains above $85\%$ for FCN and $66\%$ for AlexNet, indicating that the model tolerates a substantial fraction of client drop without significant performance degradation. In contrast, the THE-based model (where the threshold is set to $0.8$ in the experiment) shows a drastic decline in accuracy once the proportion of eliminated clients exceeds the threshold. The final model accuracy drops sharply, eventually reaching a level close to that of random guessing, indicating the failure of effective model convergence. These results demonstrate that SMHE-based model offers greater fault tolerance and flexibility in the presence of unreliable or excluded clients, whereas the THE-based system is sensitive to threshold violations due to their dependency on sufficient participant contributions.
TABLE III: Comparisons of total training time and network traffic for converged FCN and AlexNet models under different PPFL modes.

### v2

@@ 6.2 Specific Phase of the PPFL Model
  • Key generation. Each client $i$ generates its key pair $(sk_{i}=(1,s_{i}),pk_{i}=\{b_{i},a\})$ by executing $\texttt{SMHE.KeyGen}(pp)$.
@@
  • Ciphertext expandsion. The server expands the ciphertexts $ct_{i}=(c_{i,0},c_{i,1})$ $({i}\in[1,n])$ into $\overline{ct}_{i}=(\bar{c}_{i,j})_{j\in[1,n]}$ by invoking $\texttt{SMHE.Expand}(\{pk_{j}\}_{j\in[1,n]};i;ct_{i})$. Here,
  • Weighted aggregation. The server selects some clients and aggregates their local gradients as a global gradient. Denote the selected client set as $\mathcal{P}_{benign}$ with the size $m=|\mathcal{P}_{benign}|$. The server aggregates all the $m$ ciphertexts $\{\overline{ct}_{id_{i}}\}_{{id_{i}}\in\mathcal{P}_{benign}}$ as an aggregation ciphertext $\overline{ct}_{add}^{m}$. Specifically, the server computes
    $\overline{ct}_{add}^{1}$ $\leftarrow\texttt{SMHE.Add}(\alpha_{id_{1}}\overline{ct}_{id_{1}},(0)_{m+1};pk_{id_{1}},pk_{id_{m}};$
    $\{cz_{id_{1}},\bm{\Gamma}_{id_{1}}\},\{cz_{id_{m}},\bm{\Gamma}_{id_{m}}\}),\ \text{and}$
    $\overline{ct}_{add}^{i}$ $\leftarrow\texttt{SMHE.Add}(\alpha_{id_{i-1}}\overline{ct}_{add}^{i-1},\alpha_{id_{i}}\overline{ct}_{id_{i}};pk_{id_{i-1}},pk_{id_{i}};$
    $\{cz_{id_{i-1}},\bm{\Gamma}_{id_{i-1}}\},\{cz_{id_{i}},\bm{\Gamma}_{id_{i}}\})\ \text{for}\ i\in[2,m],$
    where $\alpha_{id_{i}}\in\mathbb{Q}$ serves as a possible weight in the model. Here, we have $id_{i-1}=id_{m}$ when $i=1$, and $id_{i+1}=id_{1}$ when $i=m$. The aggregation ciphertext is $\overline{ct}_{add}^{m}\overset{\triangle}{=}(\bar{c}_{0},\bar{c}_{id_{1}},\cdots,\bar{c}_{id_{m}})$ and $\bar{c}_{id_{i}}$ is then sent to the client $id_{i}$.
Partial decryption (Client side): Each client $id_{i}\in\mathcal{P}_{benign}$ computes the partial decryption result $\nu_{id_{i}}=\texttt{SMHE.PartDec}(\bar{c}_{id_{i}},$ $s_{id_{i}})$ and sends $\nu_{id_{i}}$ to the server.
Full decryption and model update (Server side):
  • Full decryption. After receiving $\{\nu_{id_{i}}\}_{id_{i}\in\mathcal{P}_{benign}}$ from all the $m$ clients, the server obtains the final decryption result $\mu\leftarrow\texttt{MKHE.Merge}(\bar{c}_{0},\{\nu_{id_{i}}\}_{id_{i}\in\mathcal{P}_{benign}})$. Then the server decodes $\mu$ as the weighted aggregation gradient $\hat{g}_{t}$.
6.3 Security
We formalize the security of the PPFL protocol instantiated with our structured SMHE from Section 5. The proof follows the standard simulation-based paradigm and relies on the RLWE assumption and the semantic security of CKKS/BFV as well as the masking mechanism in Section 5.1.
Functionality. In one aggregation round there are clients $\{P_{i}\}_{i=1}^{n}$ and a server $\mathcal{S}$. Client $P_{i}$ holds a local update $\mu_{i}$ (e.g., a gradient). The ideal functionality $\mathcal{F}_{\mathsf{Agg}}$ outputs only
$\mu\;\approx\;\sum_{i=1}^{n}\mu_{i}$
to $\mathcal{S}$ (or to all parties, as required by the application), and reveals nothing else about any individual $\mu_{i}$.
Real protocol (recap). The real system is the SMHE-based PPFL, where each client encrypts its update under CKKS augmented with a masking triple $(ct,cz,\mathbf{\Gamma})$; the server homomorphically evaluates the additive aggregation circuit; decryption is distributed: parties output partial decryptions $\nu_{i}$ over the reference set, which are linearly merged to recover $\mu$.
Adversary and leakage. We consider a semi-honest adversary $\mathcal{A}$ who may corrupt the server and an arbitrary subset $\mathcal{P}\subseteq\mathcal{P}_{benign}$ of clients while following the protocol. The leakage function $\mathcal{L}$ includes public parameters, the number of parties, the circuit/model shape, and the functionality output $\mu$.
Definition 6 (Simulation-based security target).
Let $\text{View}^{\mathrm{real}}_{\mathcal{A}}$ be the adversary’s view in the real execution (including $pp,\{pk_{i}\}$, input/expanded/evaluated ciphertexts, partial decryptions $\{\nu_{i}\}$, and optionally $\mu$). Let $\text{View}^{\mathrm{ideal}}_{\mathsf{Sim}}$ be the output of a PPT simulator given only $(pp,\mu)$. The protocol securely realizes $\mathcal{F}_{\mathsf{Agg}}$ if
$\text{View}^{\mathrm{real}}_{\mathcal{A}}(\lambda)\;\approx_{c}\;\text{View}^{\mathrm{ideal}}_{\mathsf{Sim}}(\lambda).$
Why masking matters (intuition). A CDKS-style expansion and distributed decryption enables recovering each $\mu_{i}$ from public expansion components and the corresponding partial decryption (cf. Eq. (14) in Section 4), contradicting the “only-the-sum” leakage of $\mathcal{F}_{\mathsf{Agg}}$. Our SMHE attaches a masking triple $(ct,cz,\mathbf{\Gamma})$ with an auxiliary ciphertext $c_{x}$ (derived from $\mathbf{\Gamma}$ under a second key) such that
@@ 6.3 Security
Hence, adversarially observable components are masked and become informative only upon the final linear merge.
Theorem 1 (One-round PPFL security, semi-honest).
Under the RLWE assumption and the semantic security of CKKS and the masking mechanism in Section 5.1, the SMHE-based PPFL protocol securely realizes $\mathcal{F}_{\mathsf{Agg}}$ in the semi-honest model: for any PPT adversary $\mathcal{A}$ corrupting $\mathcal{S}$ and any subset $\mathcal{P}\subseteq\mathcal{P}_{begin}$ of clients, there exists a PPT simulator $\mathsf{Sim}$ such that $\text{View}^{\mathrm{real}}_{\mathcal{A}}\approx_{c}\text{View}^{\mathrm{ideal}}_{\mathsf{Sim}}$.
Proof sketch with hybrids.
We construct $\mathsf{Sim}$ given only $(pp,\mu)$ and proceed via a sequence of hybrids.
H0 (Real execution). This is the real protocol with honest clients encrypting their $\mu_{i}$, the server evaluating, and all parties outputting partial decryptions.
H1 (Simulated honest encryptions/expansions). For each honest $i\notin\mathcal{P}$, replace input and expanded ciphertexts by fresh, distribution-consistent RLWE-type samples (same ring/modulus/layout) that are independent of $\mu_{i}$. Indistinguishability follows from the semantic security of CKKS and of the masking triple: real encryptions/masks are computationally indistinguishable from such samples.
H2 (Simulated evaluated ciphertexts). Replace all intermediate/evaluated ciphertexts derived from honest inputs by fresh RLWE-type samples with the same algebraic shape as the real evaluation. Because masking ensures that only the final linear merge un-masks the data, any single intermediate component remains computationally independent of the underlying $\mu_{i}$. Thus H1 and H2 are indistinguishable.
H3 (Simulated partial decryptions with merge consistency). For each honest party, replace $\nu_{i}$ by an RLWE-type sample drawn from the same distribution as a real partial decryption (inner product plus noise). To ensure that the final output equals $\mu$, fix one honest index $j$ and set $\tilde{\nu}_{j}$ after sampling the others so that the public linear relation used by the merge procedure satisfies
$Merge\big(\{\tilde{\nu}_{i}\},\text{public ciphertext components}\big)\approx\mu.$
This is always possible because the merge is linear and, by masking correctness,
$\langle sk,ct\rangle+\langle sk,c_{x}\rangle+\langle sk^{\prime},c_{z}\rangle\approx\mu\pmod{Q}$.
Thus H2 and H3 are indistinguishable.
Corrupted parties. For every corrupted client $i\in\mathcal{P}$, $\mathsf{Sim}$ uses the actual input and keys (already known to $\mathcal{A}$) to produce the corresponding values; for the corrupted server, $\mathsf{Sim}$ releases the simulated evaluated ciphertexts and $\{\tilde{\nu}_{i}\}$ as above.
Conclusion. Let $\text{View}^{\mathrm{ideal}}_{\mathsf{Sim}}$ be the distribution in H2. By transitivity of computational indistinguishability from H0 to H3, we obtain $\text{View}^{\mathrm{real}}_{\mathcal{A}}\approx_{c}\text{View}^{\mathrm{ideal}}_{\mathsf{Sim}}$, which proves the claim. ∎
Corollary 1 (Server + subset collusion).
Let $\mathcal{A}$ corrupt the server and any subset $\mathcal{P}$ of clients. The protocol remains simulation-secure: there exists $\mathsf{Sim}$ given $(pp,\mu)$ and $\{\mu_{i}\}_{i\in\mathcal{P}}$ (trivially known to $\mathcal{A}$) that outputs an indistinguishable view by instantiating real values for corrupted parties and applying the hybrids above to honest parties.
Multi-round composition. If each round uses fresh SMHE randomness (fresh encryption noise, fresh $(cz,\mathbf{\Gamma})$, fresh evaluations), then the per-round transcripts are independent; by a standard sequential composition theorem, over $I$ rounds the joint view is indistinguishable from an ideal system that reveals only $\{\mu^{(i)}\}_{i=1}^{I}$.
Client dropout. If some clients are absent in a round, the reference set and merge range over the active clients only. Linearity of merge and masking correctness immediately yield that only the sum over active parties is revealed; the same simulator applies unchanged.
Remark 1 (On inference from functionality leakage).
The ideal functionality $\mathcal{F}_{\mathsf{Agg}}$ reveals the aggregate $\mu$ only. However, if $\mathcal{A}$ controls n1n-1 clients, then it can in the ideal world derive the remaining honest input by subtraction: $\mu_{j}=\mu-\sum_{i\in\mathcal{P}}\mu_{i}$. Thus preventing per-client inference beyond functionality leakage requires at least two honest clients in the round.
Remark 2 (On CDKS vs. SMHE).
CDKS-style expansion plus distributed decryption admits per-client reconstruction (cf. Eq. (14)), hence cannot realize $\mathcal{F}_{\mathsf{Agg}}$ which leaks only the aggregate. In contrast, SMHE’s masking guarantees that no pre-merge component is informative about any individual $\mu_{i}$, and only the final linear merge reveals $\mu$.
6.4 Correctness
@@
We now analyze the correctness of the ciphertext aggregation and decryption processes within PPFL.
@@
During the aggregation process, the aggregated ciphertext is computed as $\overline{ct}_{\text{add}}^{1}\leftarrow\texttt{SMHE.Add}(\alpha_{id_{1}}\overline{ct}_{id_{1}},(0)_{m+1};id_{1},id_{m};$ $\{cz_{id_{1}},\bm{\Gamma}_{id_{1}}\},\{cz_{id_{m}},\bm{\Gamma}_{id_{m}}\}),$ yielding:
$\overline{ct}_{\text{add}}^{1}=\$$(\alpha_{id_{1}}c_{0}^{id_{1}}+x_{0}^{id_{1},id_{m}}+z_{0}^{id_{m}}+x_{0}^{id_{m},id_{1}}+z_{0}^{id_{1}},$
$\alpha_{id_{1}}c_{1}^{id_{1}}+x_{1}^{id_{1},id_{m}}+z_{1}^{m},0,\ldots,0,x_{1}^{id_{m},id_{1}}+z_{1}^{id_{1}}),$
@@
$\bar{c}_{0}=\sum_{id_{i}\in\mathcal{P}_{benign}}\left(\alpha_{id_{i}}c_{0}^{id_{i}}+x_{0}^{id_{i},id_{i-1}}+x_{0}^{id_{i},id_{i+1}}+2z_{0}^{id_{i}}\right),$
$\bar{c}_{id_{i}}=\alpha_{id_{i}}c_{1}^{id_{i}}+x_{1}^{id_{i},id_{i-1}}+x_{1}^{id_{i},id_{i+1}}+z_{1}^{id_{i-1}}+z_{1}^{id_{i+1}},$
for $i\in[1,m]$. Here, $cx_{id_{i},id_{j}}=(x_{0}^{id_{i},id_{j}},x_{1}^{id_{i},id_{j}})\leftarrow\texttt{Extend}(\bm{\Gamma}_{id_{i}},pk_{id_{i}},pk_{id_{j}})$ for $id_{i},id_{j}\in\mathcal{P}_{benign}$.
@@
$\qquad\mu=\bar{c}_{0}+\sum_{id_{i}\in\mathcal{P}_{benign}}(\bar{c}_{id_{i}}\cdot s_{id_{i}}+e_{id_{i}})$ (11)
$=\sum_{id_{i}\in\mathcal{P}_{benign}}\Big(\alpha_{id_{i}}\big(c_{0}^{id_{i}}+c_{1}^{id_{i}}s_{id_{i}}\big)+e_{id_{i}}+\langle sk_{id_{i}},cx_{{id_{i}},id_{i-1}}\rangle$
$\qquad+\langle sk_{id_{i}},cz_{id_{i-1}}\rangle+\langle sk_{id_{i}},cx_{id_{i},id_{i+1}}\rangle+\langle sk_{id_{i}},cz_{id_{i+1}}\rangle\Big)$
$=\alpha_{id_{i}}\mu_{id_{i}}+e^{*}\approx\sum_{id_{i}\in\mathcal{P}_{benign}}\alpha_{id_{i}}\mu_{id_{i}}\pmod{Q}$.
Eq. (11) holds due to the correctness of the masking scheme and CKKS’ scalar multiplication, which ensures that
@@
and $\langle sk_{id_{i}},\alpha_{id_{i}}ct_{id_{i}}\rangle\approx\alpha_{id_{i}}\mu_{id_{i}}\pmod{Q}$. The error bound is $\|e^{*}\|_{\infty}\leq(3m^{3}+6m^{2}+m)B_{\chi}+2dm^{2}\cdot B_{\chi}B_{\mathcal{H}}.$
@@
6.5 Implementation
6.5.1 Experimental Setup
The experiments were conducted on a computer equipped with an Nvidia GeForce GTX 1080 Ti GPU and an Intel Core i7-6700 CPU. The SMHE algorithms were implemented in C++ using the NTL 10.4.0 [31] and GMP 6.2.1 [32] libraries to handle arbitrary-length integers and high-precision arithmetic. The FL models were executed using the PyTorch 1.11.0 framework in Python 3.8. To build our PPFL framework, a dynamic library containing all SMHE-related code was generated and invoked within a Python script.
Benchmark Models, Dateset, and Parameter Setting. To evaluate the effectiveness of our SMHE-based PPFL model, we conduct comparative experiments against the existing multiparty HE schemes: CDKS [21], THE [14], and MKGSW [17]. All models are assessed under two distinct learning scenarios to ensure comprehensive and fair comparison: (i) A fully connected neural network (FCN) on the MNIST dataset. The network includes an input layer with $784$ neurons, two hidden layers with $128$ and $64$ neurons respectively, and an output layer with $10$ neurons representing digit classes from 0 to $9$ using one-hot encoding. The model is optimized using the Adam algorithm with a mini-batch size of $64$. (ii) An AlexNet network with approximately $1.25$ million parameters on the CIFAR-10 dataset. Training is conducted using a batch size of $128$ and the RMSprop optimizer with a decay factor of $10^{-6}$. For both FL tasks, the datasets are randomly partitioned among multiple clients for simulating collaborative training.
6.5.2 Experimental Results
@@
Runtime. Fig. 3 presents the runtime breakdown of PPFL models across different computation phases on both the client and server sides for a single training iteration. The results are averaged over $100$ training iterations for all models except the MKGSW-based model, whose runtime is approximated using the MKGSW evaluation time reported in TFHE [35]. As a bit-level HE scheme, MKGSW incurs an impractically high runtime due to its inherent computational complexity. For example, under the FCN setting, the total runtime reaches $3570.65$s on the client side and $2702.17$s on the server side. This excessive overhead renders MKGSW unsuitable for real-world PPFL applications. In contrast, the other three models, THE-based, CDKS-based, and our SMHE-based models, exhibit significantly lower runtimes, making them more feasible for practical deployment. Among these, THE-based and CDKS-based models demonstrate comparable runtime performance. Our SMHE-based scheme introduces a moderate increase in runtime relative to THE and CDKS, primarily due to the incorporation of the masking extension mechanism. For instance, under the AlexNet setting, the total client-side runtime of SMHE increases by less than $2\times$ compared to the CDKS- and THE-based models ($4.64$s vs. $2.72$s and $7.77$s, respectively). Similarly, on the server side, SMHE incurs a slightly higher overall runtime. In all cases, the runtime increase remains within a $2\times$ range, which we consider a reasonable trade-off given the enhanced security guarantees.
Besides, compared to THE, our approach significantly reduces the client-side key generation time (e.g., $0.08$s vs. $5.09$s for FCN) by allowing each party to generate their key locally, without the need for interactive threshold key generation. This eliminates the communication overhead inherent in THE’s distributed key setup, leading to a substantial reduction in key generation and distribution latency.
Overall, SMHE strikes a practical balance between improved security, key flexibility, and runtime efficiency. Compared to MKGSW, it achieves vastly lower computational overhead, while maintaining acceptable runtime increases relative to THE and CDKS. These features make SMHE a scalable and effective solution for MPC applications.
TABLE II: Communication traffic comparison.
@@
Communication Traffic. Table II presents the communication traffic per iteration for different PPFL models under FCN and AlexNet networks. The row labeled “Plain (#gradient)” reflects the number of model gradients in the plain FL, 109,386109{,}386K for FCN and 1.251.25M for AlexNet. For encrypted models, the MKGSW-based model introduces an extremely high communication burden, 33.3933.39GB per iteration for FCN and 381.47381.47GB for AlexNet, mainly due to its bit-level encryption. This renders MKGSW impractical for real-world PPFL applications, where communication efficiency is a critical concern. In contrast, the other three models: THE-, CDKS-, and SMHE-based scheme offer significantly lower communication costs and are more suitable for practical deployment. Specifically, THE- and CDKS-based models share similar ciphertext formats and therefore exhibit identical communication traffic, requiring only 77MB and 76.576.5MB per iteration for FCN and AlexNet, respectively. Our SMHE-based model incurs a moderate increase in communication traffic, reaching 1616MB for FCN and 139.5139.5MB for AlexNet. This overhead stems from the ciphertext extension mechanism introduced by our masking design. However, despite the increase, the total traffic remains within approximately $2\times$ that of CDKS and THE, which is a reasonable cost given the enhanced security and flexibility.
Model Accuracy. Fig. 4 displays the final model accuracy of the PPFL models using SMHE, CDKS, and THE under different client elimination rates. As observed from the results, the SMHE (CDKS)-based model exhibits strong robustness against client elimination [^3] While SMHE requires all participants for decryption, its full key independence and dynamism make it particularly suitable for cross-silo FL settings, where participants may join or leave over time but typically remain reliable within each aggregation session.. Even with $50$ distributed clients and a high elimination rate of $50\%$, the final model accuracy remains above $85\%$ for FCN and $66\%$ for AlexNet, indicating that the model tolerates a substantial fraction of client drop without significant performance degradation. In contrast, the THE-based model (where the threshold is set to $0.8$ in the experiment) shows a drastic decline in accuracy once the proportion of eliminated clients exceeds the threshold. The final model accuracy drops sharply, eventually reaching a level close to that of random guessing, indicating the failure of effective model convergence. These results demonstrate that SMHE-based model offers greater fault tolerance and flexibility in the presence of unreliable or excluded clients, whereas the THE-based system is sensitive to threshold violations due to their dependency on sufficient participant contributions.
TABLE III: Comparisons of total training time and network traffic for converged FCN and AlexNet networks under different PPFL modes.